# FRABench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

### v2

@@ C.3 Aspects Used in the benchmark
Table 29: Aspects used by evaluators in the public benchmark evaluations. (Part 3)
Appendix D More Experiment Results
D.1 Specific Experimental Results
@@
We report the specific average accuracy of each evaluator for every aspect within each task in Figures 4 and 6. The alignment is measured against: (1) GPT-4o on the FRA-ID and FRA-OOD, and (2) human annotators on the FRA-ID-H and FRA-OOD-H. Detailed results are presented in Table 30 to 41. The reported average accuracy is computed by aggregating the same aspect across different sub-tasks within the same task. We also include the comprehensive results on MLLM-as-a-Judge in Table 50. The experimental results of FRAUAs-OOD are presented in Figure 13 and Table 46 to 49.
@@
Action Consistency ActC 88.2 64.9 45.2 24.5 84.7
Bias Bias 87.4 74.7 67.6 46.2 80.4
Character Consistency ChaC 79.1 58.2 33.7 14.2 72.8
Coherence Coh 95.0 93.7 85.3 55.1 94.5
Completeness Com 86.0 83.2 77.7 38.2 87.6
Completeness Com 90.5 88.0 80.1 46.0 88.8
Creativity Cre 90.8 88.0 76.6 35.1 86.5
Engagingness Eng 88.8 88.7 78.8 29.6 90.0
Feasibility Fea 90.7 87.3 82.5 50.2 91.4
Fluency Flu 98.2 93.1 85.8 49.5 94.6
Helpfulness Hel 94.7 90.3 79.6 46.6 90.3
Image Repetition ImgR 86.5 71.0 43.1 37.3 81.7
Readability Rea 94.0 90.4 81.5 53.1 92.4
Safety Saf 84.5 79.5 53.0 31.3 76.1
Scene Consistency SceC 90.5 78.6 35.0 28.6 88.1
Semantic Consistency SemC 93.4 73.6 50.0 28.7 90.6
Stylistic Consistency StyC 90.0 66.1 32.7 20.8 83.5
Toxicity Tox 97.5 84.2 87.8 54.1 93.7
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 50.0 43.8 28.1 34.4 51.0
Alignment Ali 67.5 70.0 47.5 45.0 55.0
Attribute Alignment AttA 62.5 62.5 52.5 37.5 65.0
Color Alignment ColA 65.0 60.0 57.5 37.5 49.5
Count Alignment CouA 55.0 60.0 50.0 20.0 60.0
Fidelity Fid 30.0 42.5 40.0 33.0 35.0
Harmfulness Har 100 62.5 95.0 82.5 95.0
Object Alignment ObjA 52.5 57.5 50.0 27.5 57.5
Object Relationship ObjR 60.0 50.0 45.0 32.5 52.5
Shape Alignment ShaA 65.0 55.0 60.0 35.0 70.0
Spatial Alignment SpaA 57.5 60.0 37.5 27.5 57.5
Texture Alignment TexA 52.5 60.0 50.0 50.0 62.5
@@
Action Consistency ActC 52.0 44.0 33.3 33.3 53.3
Bias Bias 67.5 57.5 65.0 37.5 67.5
Character Consistency ChaC 50.0 30.0 37.5 32.5 38.5
Clarity Cla 87.5 87.5 82.5 55.0 90.0
Coherence Coh 73.8 85.0 78.7 41.2 85.0
Completeness Com 70.0 67.5 72.5 30.0 72.5
Completeness Com 77.8 80.6 83.3 38.9 80.6
Creativity Cre 62.5 62.5 47.5 35.0 67.5
Engagingness Eng 80.0 82.5 75.0 32.5 77.5
Feasibility Fea 86.1 91.7 83.3 44.4 91.7
Fluency Flu 75.0 76.2 75.0 46.3 76.0
Grammaticality Gra 62.5 60.0 62.5 41.2 64.0
Helpfulness Hel 80.6 83.3 75.0 38.9 77.8
Image Repetition ImgR 61.3 42.7 41.3 36.0 66.7
Readability Rea 65.0 67.5 62.5 38.8 67.5
Safety Saf 58.3 63.9 55.6 33.3 50.0
Scene Consistency SceC 64.0 61.3 33.3 33.3 62.7
Semantic Consistency SemC 53.3 48.0 44.0 28.0 56.0
Stylistic Consistency StyC 65.3 45.3 32.0 36.0 53.3
Toxicity Tox 82.5 90.0 92.5 37.5 90.0
@@
Action Consistency ActC 90.0 62.5 32.2 42.2 71.1
Character Consistency ChaC 85.6 71.3 20.0 24.4 81.1
Clarity Cla 87.5 87.5 80.0 50.0 92.5
Coherence Coh 92.2 95.0 92.2 68.9 96.7
Fluency Flu 92.2 87.5 87.8 60.0 88.9
Harmfulness Har 91.1 65.0 81.1 61.1 76.7
Image Coherence ImaC 90.0 67.1 46.7 30.0 81.1
Image Repetition ImgR 84.4 75.0 46.7 46.7 75.6
Readability Rea 96.7 86.3 88.9 57.8 88.9
Scene Consistency SceC 85.6 76.2 53.3 31.1 86.7
Semantic Consistency SemC 74.4 73.4 53.3 32.2 84.4
Stylistic Consistency StyC 88.9 61.3 37.8 16.7 73.3
Fulfillment Full 90.0 95.0 85.0 42.5 95.0
Originality Orig 84.0 70.0 74.0 50.0 80.0
Reasonableness Reas 100 97.5 92.5 30.0 92.5
Simplification Simp 90.0 77.5 84.0 52.0 90.0
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
@@
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Consistency ActC 77.5 63.4 25.0 50.0 67.5
Character Consistency ChaC 60.0 68.7 22.5 35.0 68.8
Clarity Cla 82.5 90.0 90.0 47.5 87.5
Coherence Coh 86.3 95.0 92.5 62.5 96.3
Fluency Flu 90.0 95.0 97.5 65.0 95.0
Harmfulness Har 73.8 71.3 67.5 65.0 72.5
Image Coherence ImaC 87.0 67.1 46.7 30.0 81.1
Image Repetition ImgR 70.0 53.6 47.5 45.0 71.3
Readability Rea 85.0 90.0 91.2 60.0 91.2
Scene Consistency SceC 76.2 85.5 58.8 38.8 81.2
Semantic Consistency SemC 74.4 73.4 53.3 32.2 84.4
Stylistic Consistency StyC 62.5 75.7 40.0 20.0 68.8
Fulfillment Full 95.0 97.3 82.5 40.0 97.5
Originality Orig 92.5 92.9 77.5 55.0 85.0
Reasonableness Reas 95.0 94.7 85.0 32.5 90.0
Simplification Simp 90.0 79.3 90.0 50.0 90.0
Figure 13: Comparison on the FRAUAs-OOD to evaluate the generalization in UAs.
Method GPT-4o (FRAUAs-OOD)
NLG IU ITIG
GPT-4o 89.2 86.3 90.6
Claude-3.5 77.3 77.6 83.0
Qwen2VL-72B 69.7 71.4 65.3
Qwen2VL-7B 49.9 56.5 44.4
Themis 54.2 - -
LLaVA-Critic - 50.0 -
Ours 82.8 83.1 90.4
Table 46: Comparison on FRAUAs-OOD. We compute the average accuracy of UAs for each task.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Themis Ours
Clarity Cla 90.0 82.5 65.8 42.5 36.7 84.2
Coherence Coh 89.2 90.8 78.3 50.0 40.8 83.3
Fluency Flu 91.7 89.2 71.7 49.2 56.5 81.7
Grammaticality Gra 86.7 51.0 65.8 50.0 68.8 82.9
Harmfulness Har 92.4 81.0 90.9 65.3 80.6 93.9
Readability Rea 89.4 86.7 72.0 48.0 38.6 85.5
Sentence Complexity SenC 89.7 75.3 60.8 52.7 38.5 82.0
Spelling Accuracy SpeA 91.2 84.1 78.0 38.9 78.9 86.3
Structure Accuracy StrA 82.4 74.4 63.4 53.3 52.4 80.2
Tense Consistency TenC 96.2 61.5 80.8 64.0 73.1 80.8
Text Repetition TexR 90.0 78.5 52.0 38.2 41.5 77.5
Vocabulary Complexity VocC 81.7 72.6 57.5 47.6 44.7 75.5
Table 47: The specific experimental results of the NLG task in the FRAUAs-OOD are used to evaluate generalization in UAs.
Bias Bias 88.3 86.4 82.6 64.3 23.8 86.2
Clarity Cla 75.2 63.5 53.7 51.0 49.5 66.7
Coherence Coh 91.8 87.3 73.3 50.3 73.8 89.7
Ethnic Bias EthB 99.5 98.5 98.0 83.9 14.7 99.8
Fluency Flu 81.6 76.0 71.3 52.6 67.2 80.0
Grammaticality Gra 90.5 56.8 83.0 65.0 37.7 89.7
Readability Rea 88.8 84.0 70.6 50.7 72.6 85.2
Sentence Complexity SenC 86.8 74.3 47.0 43.6 48.4 86.0
Vocabulary Complexity VocC 74.6 71.7 63.1 47.7 62.6 65.5
Table 48: The specific experimental results of the IU task in the FRAUAs-OOD are used to evaluate generalization in UAs.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Clarity Cla 94.2 88.6 79.4 57.2 92.8
Coherence Coh 93.8 90.5 84.2 55.4 92.9
Fluency Flu 97.1 91.3 84.5 50.6 95.3
Grammaticality Gra 92.6 85.1 76.2 42.1 88.0
Image Repetition ImgR 82.5 71.6 39.9 38.5 84.7
Readability Rea 95.3 89.6 83.8 51.8 95.1
Scene Consistency SceC 83.9 73.4 34.8 30.8 89.4
Semantic Consistency SemC 85.9 73.8 40.0 29.6 85.8
Table 49: The specific experimental results of the ITIG task in the FRAUAs-OOD are used to evaluate generalization in UAs.
Settings MLLM COCO C.C. Diff Graphics Math Text WIT Chart VisIT CC-3M M2W SciQA Aes MM-Vet Ave.
tau($\uparrow$)↑ ) GPT-4V* 69.6 82.4 84.7 63.9 56.4 67.3 67.9 65.7 64.0 61.2 52.1 41.5 60.6 52.9 63.6
@@
Table 50: Comprehensive results on MLLM-as-a-Judge. *: The results for GPT-4V and LLaVA-Critic are taken from the original papers [29] and [12], respectively. All other results are obtained from our experiments using their publicly available codebases.
D.2 Staged Training Experiment Results
During GenEval training, we compared unified training and staged training across tasks. For staged training, we group tasks with similar output modalities: NLG and IU as Text, IG as Image, and ITIG as Inter. Using the fine-tuning method from Appendix C.1, we train models in stages, swapping the order of these task types, resulting in six models. Experiments on the FRA-ID and FRA-ID-H (Figure 14 and Table 51) show that the model performs best on tasks from the last-trained category. In contrast, unified training provided more balanced performance across tasks.
Figure 14: Experimental results in the FRA-ID and FRA-ID-H using different types of training set orders for staged training.
Method GPT-4o Human
NLG IU IG ITIG NLG IU IG ITIG
Image-Inter-Text 89.2 87.2 63.9 85.0 77.8 75.2 56.5 68.2
Image-Text-Inter 86.6 84.7 61.8 87.1 76.9 73.1 55.1 69.4
Inter-Image-Text 89.2 86.9 66.0 82.1 78.4 75.5 54.3 67.4
Inter-Text-Image 85.2 83.1 71.5 81.2 76.1 74.1 58.2 65.4
Text-Image-Inter 84.2 83.7 65.1 87.2 75.1 72.7 58.4 70.2
Text-Inter-Image 84.6 81.5 70.3 83.3 75.7 74.3 58.5 67.5
Ours 89.0 86.3 70.0 87.0 80.0 74.4 59.2 69.4
Table 51: Average accuracy for staged training results on FRA-ID and FRA-ID-H. Unified training obtains balanced or better performance.
D.3 GenEval-72B Experiment Results
We also train GenEval-72B using the same fine-tuning method as GenEval-7B with Qwen2-VL-72B-Instruct as the backbone and test it under the same experiment settings. The results, shown in Figures 15 and 16 and Table 52 to 54, indicate that GenEval-72B offers some improvement over GenEval-7B. However, due to performance and deployment trade-offs, we find GenEval-7B to be more cost-effective.
Figure 15: Comparison GenEval-72B on FRA-ID and FRA-ID-H. We computed the average accuracy for both UAs and TAs across GPT-4o Alignment and Human Alignment conditions.
Method GenAI-Bench Winoground
tau($\uparrow$) diff($\uparrow$) Relation Object Both Ave.
@@ D.3 GenEval-72B Experiment Results
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4
Ours 53.6 65.5 57.5 59.1 80.7 65.7
Ours-72B 55.5 69.0 61.5 63.4 62.5 62.4
Table 52: The results of GenEval-72B compared with different models on two IG benchmarks.
Figure 16: Comparison GenEval-72B on FRA-OOD and FRA-OOD-H. “ow/unseen” denotes the average accuracy calculated only for unseen TAs. The aspect marked in blue is unseen TAs.
@@
Table 53: The results of GenEval-72B compared with different models on three NLG benchmarks.
@@
Table 54: The results of GenEval-72B compared with different models on three IU benchmarks.
Appendix E Experimental Details for Human Evaluation
The user interface used for human evaluation is shown in Figures 17 and 18. For each aspect, we adopt a 5-point Likert scale. Additionally, to account for cases where certain instructions are not applicable to TAs, we include a "-1" option to indicate inapplicability. To maintain high-quality annotation, we engage three experts to annotate each data point concurrently. These experts, recruited from graduate students in relevant fields to enhance annotation reliability, receive specialized training that goes beyond assessing response correctness, enabling them to focus on multiple aspects such as helpfulness, clarity, comprehensiveness, and more.
We also compute the Inter-Annotator Agreement (IAA) for each annotator, with detailed results presented in Table 55 and Table 56. The final IAA among the three annotators achieved average scores of 84.2, 87.4, and 93.8 for FRA-ID-H, and 84.0, 81.2, 78.2 for FRA-ODD-H, respectively. These results demonstrate a high level of reliability in our test dataset.
Annotator NLG IU IG ITIG
@@ Appendix E Experimental Details for Human Evaluation
Appendix F The Prompt Template Used by UAs and TAs
In this section, we provide the prompt template used for training/inferencing GenEval in Table 57 to Table 64. Note that We employed distinct prompt templates for UAs and TAs to enable specific aspect evaluation while avoiding overall assessment.
[BEGIN DATA]
[Criterion]: {criterion_description}
[Response 1]: {response_1}
[Response 2]: {response_2}
[END DATA]
Here are the instructions to assess and compare the two responses:
Table 57: The example prompt templates for UAs only include the text output modality.
[BEGIN DATA]
[Criterion]: {criterion_description}
[END DATA]
Here are the instructions to assess and compare the two images:
Table 58: The example prompt templates for UAs only include the image output modality.
@@ Appendix F The Prompt Template Used by UAs and TAs
Table 59: The example prompt templates for UAs only include the multi-image output.
@@
Table 60: The example prompt templates for TAs in NLG task.
@@
Table 61: The example prompt templates for TAs in IU task.

### v3

@@ Appendix C Appendix C: Experiment Detail
Table 27: Aspects used by evaluators in the meta-benchmark evaluations. (Part 4)
Appendix D Appendix D: More Experiment Results
In-domain Evaluation
For in-domain evaluation, we assess performance using the FRA-ID and FRA-ID-H. Figure 11 presents radar charts that illustrate the overall degree of alignment. The accompanying Table 29 reports average accuracies for each task. Additionally, we also sample three types of aspects for aspect-level analysis. As shown in Figure 12, we select aspects from four test sets based on UFEval’s performance: (1) Weak aspects, where the best model surpasses UFEval by over 20%. (2) Universal poor aspects, where scores from all evaluators are very low. (3) Strong aspects, where UFEval excels. Only two weak aspects appear across all test sets, and both are included. For the other categories, we choose representative subsets.
The results show that UFEval achieves higher alignment with both GPT-4o and human annotators compared to the baselines (see shaded area in radar chart). Specifically, excluding GPT-4o, UFEval surpasses all baselines on NLG, IU, and ITIG tasks within FRA-ID and FRA-ID-H, achieving overall average accuracies of 82.4% and 72.1%, respectively. On NLG, UFEval outperforms Themis, while our base model Qwen2Vl-7B performs much worse. Moreover, UFEval shows strong performance on the ITIG of FRA-ID-H, where its alignment with human annotators exceeds that of GPT-4o. This improvement can be attributed to UFEval’s generalization ability and cross-task transfer performance.
Experimental Results for FRAUAs-OOD
In the training set, for each sub-task, according to our aspect selection procedure, if the task output modality includes text, all aspects under the Text Branch in the UAs Tree can be selected. However, once the model learns a UAs in one sub-task, it can leverage this UAs for inference in other sub-tasks without requiring retraining. Additionally, to maintain sample balance between UAs and TAs during training (as selecting all aspects under the Text Branch for a text-output task would lead to an imbalance in the number of UAs and TAs in the training set), we selectively sample a subset of corresponding UAs in the training set.
To validate that UFEval can evaluate UAs in sub-tasks where these UAs were not encountered during training, we construct FRAUAs-OOD, which comprises the same sub-tasks as the training set but introduces unseen UAs across 4 tasks. The experimental results are shown in Figure 13 and Table 28. The figure illustrates that UFEval exhibits robust generalization capability despite not being trained on the complete set of UAs for each sub-task.
Specific Experimental Results
@@ Appendix D Appendix D: More Experiment Results
We report the specific average accuracy of each evaluator for every aspect within each task in Figure 11 and Figure 3 of the main text. The alignment is measured against: (1) GPT-4o on the FRA-ID and FRA-OOD, and (2) human annotators on the FRA-ID-H and FRA-OOD-H. Detailed results are presented in Tables 30 to 37. The reported average accuracy is computed by aggregating the same aspect across different sub-tasks within the same task. We also include the comprehensive results on MLLM-as-a-Judge in Table 46. The specific accuracy of multi-aspect assessment learning is show in Table 52 and  54.
Experiment Results of UFEval-72B
We also train UFEval-72B using the same fine-tuning method as UFEval-7B with Qwen2-VL-72B-Instruct as the backbone and test it under the same experiment settings. The results, shown in Figures 14 and 15 and Tables 47 to 49, indicate that UFEval-72B offers some improvement over UFEval-7B.
Appendix E Appendix E: Human Annotation
The user interface used for human annotation is shown in Figures 16 and 17. For each aspect, we adopt a 5-point Likert scale. Additionally, to account for cases where certain instructions are not applicable to TAs, we include a "-1" option to indicate inapplicability. To maintain high-quality annotation, we engage three experts to annotate each data point concurrently. These experts, recruited from graduate students to enhance annotation reliability, are provided with guidance whenever they encounter challenges during the annotation process, enabling them to focus on multiple aspects such as helpfulness, clarity, comprehensiveness, and more.
We also compute the Inter-Annotator Agreement (IAA) for each annotator, with detailed results presented in Table 50 and Table 51. The final IAA among the three annotators achieved average scores of 84.2, 87.4, and 90.8 for FRA-ID-H, and 84.0, 81.2, 78.2 for FRA-ODD-H, respectively. These results demonstrate a high level of reliability in our test dataset.
Appendix F Appendix F: The Prompt Template
In this section, we provide the prompt template used for GPT-4o annotation and training/inferencing UFEval. Note that we employed distinct prompt templates for UAs and TAs to enable specific aspect evaluation while avoiding overall assessment.
Appendix G Appendix G: Qualitative Examples
Table 64, Table 65, Table 66, Tables 67 to 69 show some good qualitative examples of feedback generated by UFEval. Table 70, Table 71, and Table 72 shows a relatively inferior example of feedback generated by UFEval.
Method GPT-4o Alignment
NLG IU ITIG
Claude-3.5 77.3 77.6 83.0
Qwen2VL-72B 69.7 71.4 65.3
Qwen2VL-7B 49.9 56.5 44.4
Themis 54.2 - -
LLaVA-Critic - 50.0 -
Ours 82.8 83.1 90.4
Table 28: Average accuracy on FRAUAs-OOD.
Method Seen UAs for Seen Tasks Seen TAs for Seen Tasks
FRA-ID (GPT-4o) FRA-ID-H (Human) FRA-ID (GPT-4o) FRA-ID-H (Human)
NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG
GPT-4o - - - - 76.1 79.1 65.0 66.1 - - - - 79.5 74.8 58.8 73.6
Themis 60.4 - - - 60.3 - - - 47.5 - - - 53.4 - - -
LLaVA-Critic - 47.5 - - - 39.0 - - - 68.0 - - - 56.0 - -
Figure 13: Comparison on FRAUAs-OOD to evaluate the generalization in UAs.
@@ Appendix G Appendix G: Qualitative Examples
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Consistency ActC 88.2 64.9 45.2 24.5 84.7
Bias Bias 87.4 74.7 67.6 46.2 80.4
Character Consistency ChaC 79.1 58.2 33.7 14.2 72.8
Coherence Coh 95.0 93.7 85.3 55.1 94.5
Completeness Com 86.0 83.2 77.7 38.2 87.6
Completeness Com 90.5 88.0 80.1 46.0 88.8
Creativity Cre 90.8 88.0 76.6 35.1 86.5
Engagingness Eng 88.8 88.7 78.8 29.6 90.0
Feasibility Fea 90.7 87.3 82.5 50.2 91.4
Fluency Flu 98.2 93.1 85.8 49.5 94.6
Helpfulness Hel 94.7 90.3 79.6 46.6 90.3
Image Repetition ImgR 86.5 71.0 43.1 37.3 81.7
Readability Rea 94.0 90.4 81.5 53.1 92.4
Safety Saf 84.5 79.5 53.0 31.3 76.1
Scene Consistency SceC 90.5 78.6 35.0 28.6 88.1
Semantic Consistency SemC 93.4 73.6 50.0 28.7 90.6
Stylistic Consistency StyC 90.0 66.1 32.7 20.8 83.5
Toxicity Tox 97.5 84.2 87.8 54.1 93.7
@@
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 50.0 43.8 28.1 34.4 51.0
Alignment Ali 67.5 70.0 47.5 45.0 55.0
Attribute Alignment AttA 62.5 62.5 52.5 37.5 65.0
Color Alignment ColA 65.0 60.0 57.5 37.5 49.5
Count Alignment CouA 55.0 60.0 50.0 20.0 60.0
Fidelity Fid 30.0 42.5 40.0 33.0 35.0
Harmfulness Har 100 62.5 95.0 82.5 95.0
Object Alignment ObjA 52.5 57.5 50.0 27.5 57.5
Object Relationship ObjR 60.0 50.0 45.0 32.5 52.5
Shape Alignment ShaA 65.0 55.0 60.0 35.0 70.0
Spatial Alignment SpaA 57.5 60.0 37.5 27.5 57.5
Texture Alignment TexA 52.5 60.0 50.0 50.0 62.5
@@
Action Consistency ActC 52.0 44.0 33.3 33.3 53.3
Bias Bias 67.5 57.5 65.0 37.5 67.5
Character Consistency ChaC 50.0 30.0 37.5 32.5 38.5
Clarity Cla 87.5 87.5 82.5 55.0 90.0
Coherence Coh 73.8 85.0 78.7 41.2 85.0
Completeness Com 70.0 67.5 72.5 30.0 72.5
Completeness Com 77.8 80.6 83.3 38.9 80.6
Creativity Cre 62.5 62.5 47.5 35.0 67.5
Engagingness Eng 80.0 82.5 75.0 32.5 77.5
Feasibility Fea 86.1 91.7 83.3 44.4 91.7
Fluency Flu 75.0 76.2 75.0 46.3 76.0
Grammaticality Gra 62.5 60.0 62.5 41.2 64.0
Helpfulness Hel 80.6 83.3 75.0 38.9 77.8
Image Repetition ImgR 61.3 42.7 41.3 36.0 66.7
Readability Rea 65.0 67.5 62.5 38.8 67.5
Safety Saf 58.3 63.9 55.6 33.3 50.0
Scene Consistency SceC 64.0 61.3 33.3 33.3 62.7
Semantic Consistency SemC 53.3 48.0 44.0 28.0 56.0
Stylistic Consistency StyC 65.3 45.3 32.0 36.0 53.3
Toxicity Tox 82.5 90.0 92.5 37.5 90.0
@@
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Consistency ActC 90.0 62.5 32.2 42.2 71.1
Character Consistency ChaC 85.6 71.3 20.0 24.4 81.1
Clarity Cla 87.5 87.5 80.0 50.0 92.5
Coherence Coh 92.2 95.0 92.2 68.9 96.7
Fluency Flu 92.2 87.5 87.8 60.0 88.9
Harmfulness Har 91.1 65.0 81.1 61.1 76.7
Image Coherence ImaC 90.0 67.1 46.7 30.0 81.1
Image Repetition ImgR 84.4 75.0 46.7 46.7 75.6
Readability Rea 96.7 86.3 88.9 57.8 88.9
Scene Consistency SceC 85.6 76.2 53.3 31.1 86.7
Semantic Consistency SemC 74.4 73.4 53.3 32.2 84.4
Stylistic Consistency StyC 88.9 61.3 37.8 16.7 73.3
Fulfillment Full 90.0 95.0 85.0 42.5 95.0
Originality Orig 84.0 70.0 74.0 50.0 80.0
Reasonableness Reas 100 97.5 92.5 30.0 92.5
Simplification Simp 90.0 77.5 84.0 52.0 90.0
@@
Action Consistency ActC 77.5 63.4 25.0 50.0 67.5
Character Consistency ChaC 60.0 68.7 22.5 35.0 68.8
Clarity Cla 82.5 90.0 90.0 47.5 87.5
Coherence Coh 86.3 95.0 92.5 62.5 96.3
Fluency Flu 90.0 95.0 97.5 65.0 95.0
Harmfulness Har 73.8 71.3 67.5 65.0 72.5
Image Coherence ImaC 87.0 67.1 46.7 30.0 81.1
Image Repetition ImgR 70.0 53.6 47.5 45.0 71.3
Readability Rea 85.0 90.0 91.2 60.0 91.2
Scene Consistency SceC 76.2 85.5 58.8 38.8 81.2
Semantic Consistency SemC 74.4 73.4 53.3 32.2 84.4
Stylistic Consistency StyC 62.5 75.7 40.0 20.0 68.8
Fulfillment Full 95.0 97.3 82.5 40.0 97.5
Originality Orig 92.5 92.9 77.5 55.0 85.0
Reasonableness Reas 95.0 94.7 85.0 32.5 90.0
Simplification Simp 90.0 79.3 90.0 50.0 90.0
@@
Settings MLLM COCO $C.$ Diff Graphics Math Text WIT Chart VisIT CC-3M M2W SciQA Aes $M$-Vet Ave.
tau($\uparrow)$ GPT-4V* 69.6 82.4 84.7 63.9 56.4 67.3 67.9 65.7 64.0 61.2 52.1 41.5 60.6 52.9 63.6
@@
Table 46: Comprehensive results on MLLM-as-a-Judge. *: The results for GPT-4V and LLaVA-Critic are taken from the original papers (Chen et al. 2024) and (Xiong et al. 2024), respectively. All other results are obtained from our experiments using their publicly available codebases.
Figure 14: Comparison of UFEval-72B on FRA-ID and FRA-ID-H. The top shows alignment with GPT-4o using FRA-ID, the bottom shows alignment with human annotators using FRA-ID-H. Each point is the average accuracy (with ties) for an aspect shared across sub-tasks of the same task.
Method GenAI-Bench Winoground Pickapic
tau($\uparrow$) diff($\uparrow$) Relation Object Both Ave. tau($\uparrow$) diff($\uparrow$)
@@
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5 54.4 59.2
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1 49.1 53.7
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0 38.6 38.3
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6 38.1 40.2
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4 48.9 58.0
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4 48.8 55.7
Ours 53.6 65.5 57.5 59.1 80.7 65.7 50.0 57.3
Ours-72B 55.5 69.0 61.5 63.4 62.5 62.4 52.0 58.6
Table 47: Evaluation as LLM-as-a-Judge for IG task. We evaluate baselines across three meta-benchmarks.
@@
Table 48: Evaluation as LLM-as-a-Judge for NLG task. We evaluate baselines across three meta-benchmarks.
@@
Table 49: Evaluation as LLM-as-a-Judge for IU tasks. We evaluate baselines across three meta-benchmarks.
Figure 15: Comparison of UFEval-72B on FRA-OOD and FRA-OOD-H. The blue-colored aspects indicate unseen TAs, whereas the black-colored aspects represent seen UAs. The ’ow/ unseen’ designation in the table represents evaluations conducted exclusively on unseen aspects, with accuracy metrics computed only for unseen TAs.
Annotators NLG IU IG ITIG
@@
Aspect Abbr Qwen2VL-7B w/ IU w/ IU+ITIG w/ IU+IG Ours
Accuracy Acc 54.5 80.1 83.1 85.8 85.5
Action Alignment ActA 37.4 65.0 69.8 69.3 69.6
Alignment Ali 42.2 75.2 77.6 80.0 80.0
Attribute Alignment AttA 51.7 77.2 79.0 81.5 84.5
Color Alignment ColA 41.5 74.0 78.5 84.5 85.5
Completeness Com 61.3 90.2 94.5 94.3 95.0
Conciseness Con 31.8 94.5 94.5 95.3 95.0
Context Inconsistency ConI 44.3 80.7 84.0 83.5 84.3
Count Alignment CouA 50.7 75.0 75.7 76.5 80.0
Helpfulness Hel 52.5 90.7 91.2 93.0 94.8
Instruction Following InsF 48.4 81.7 82.7 84.1 83.7
Instruction Inconsistency InsI 42.9 76.7 79.2 84.0 85.5
Object Alignment ObjA 46.1 68.8 75.5 77.5 78.1
Shape Alignment ShaA 49.7 76.2 77.5 80.0 81.0
Spatial Alignment SpaA 53.7 76.0 81.5 86.8 86.8
Text-Image Relationship T-IR 42.0 86.3 87.5 91.5 92.5
Texture Alignment TexA 47.2 78.0 80.0 85.5 85.8
Table 52: The specific experimental results of multi-aspect assessment learning in the IU tasks.
Aspect Abbr Qwen2VL-7B w/ IG w/ IG+ITIG IG+IU Ours
Accuracy Acc 35.9 67.8 68.4 77.2 77.0
Action Alignment ActA 48.8 70.9 71.5 72.8 73.8
Alignment Ali 39.1 63.5 65.6 70.6 69.7
Attribute Alignment AttA 40.7 64.9 66.7 71.0 71.3
Color Alignment ColA 40.7 68.1 70.3 74.6 76.6
Count Alignment CouA 52.1 74.3 74.9 80.1 80.7
Fidelity Fid 40.8 44.0 42.0 44.0 44.0
Harmfulness Har 76.1 76.5 78.7 77.8 77.7
Object Alignment ObjA 53.5 65.3 70.0 71.7 73.3
Object Relationship ObjR 38.9 63.0 64.7 68.1 70.4
Shape Alignment ShaA 32.0 54.1 58.2 60.2 62.2
Spatial Alignment SpaA 50.4 66.0 69.2 70.4 73.2
Texture Alignment TexA 37.4 56.7 58.4 60.2 61.3
Table 53: The specific experimental results of multi-aspect assessment learning in the IG tasks.
Aspect Abbr Qwen2VL-7B w/ ITIG w/ ITIG+IU w/ ITIG+IG Ours
Action Consistency ActC 24.5 71.1 75.5 83.7 84.7
Character Consistency ChaC 14.2 64.1 67.1 70.4 72.8
Completeness Com 38.2 87.0 87.6 87.8 87.6
Completeness Com 46.0 87.0 88.5 89.0 88.8
Creativity Cre 35.1 83.3 83.3 85.2 86.5
Engagingness Eng 29.6 86.3 86.9 89.0 90.0
Feasibility Fea 50.2 86.9 91.4 90.6 91.4
Helpfulness Hel 46.6 87.4 88.3 90.6 90.3
Image Repetition ImgR 37.3 73.9 75.0 80.5 81.7
Safety Saf 31.3 69.4 71.8 72.8 76.1
Scene Consistency SceC 28.6 77.2 79.9 85.6 88.1
Semantic Consistency SemC 28.7 80.4 85.6 89.7 90.6
Stylistic Consistency StyC 20.8 75.7 80.1 82.0 83.5
Table 54: The specific experimental results of multi-aspect assessment learning in the ITIG tasks.
@@
Table 55: The prompt templates for Universal Aspects only include the text-and-image output modality.
@@
Table 56: The prompt templates for Task-specific Aspects in NLG task.
@@
Table 57: The prompt templates for Task-specific Aspects in IU task.